OpenStudy (anonymous):

erin and susan both took their dogs for a walk. Together theyw alked 15 miles. If erin walked 3 miles less than susan how far did susan walk with her dog?

OpenStudy (mertsj):

e=s-3 e+s=15 Replace e in the second equation wis s-3 from the first equation and solve.

OpenStudy (anonymous):

Let distance covered by susan = x miles then distance covered by erin = (x - 3) miles together they walked 15 miles menas x + (x - 3) = 15 2x - 3 = 15 2x = 15 + 3 x = 18/2 x = 9 so susan walked 9 miles and erin walked (9-3) = 6 miles